OpenStudy (anonymous):

What is the magnitude and direction of if G(3, -2) and H(6, 4)? magnitude: 6.71 units; direction 63.43° magnitude: 6.71 units: direction: 26.57° magnitude: 3.61 units; direction: 26.57° magnitude: 3.61 units: direction: 63.43°

ganeshie8 (ganeshie8):

magnitude is just distance between the two points, use the formula : \[\sqrt{(x2-x1)^2 + (y2-y1)^2}\] direction formula : \[\theta = tan^-1\ \frac{y2-y2}{x2-x1}\]

ganeshie8 (ganeshie8):

its \[\theta = tan^-1\ \frac{y2-y1}{x2-x1} \]